fromplanks law deduce the value of frequency corresponding to peak of (E(not)- V) curve at temp 1000k. In what spectral region does this frequency lie.
1
Expert's answer
2020-03-26T12:16:43-0400

The low of Plank


"\epsilon(\nu,T)=\frac{2h\nu^3}{c^2}\cdot\frac{1}{e^{\frac{h\nu}{kT}}-1}."


"\frac{d\epsilon(\nu,T)}{d\nu}=0"


So, we get that


"-\frac{h\nu}{kT}\cdot\frac{e^{\frac{h\nu}{kT}}}{e^{\frac{h\nu}{kT}}-1}+3=0" "\to" "(x-3)e^x+3=0" "(x=\frac{h\nu}{kT})"


From this equation "x = 2.821439372122"


"\nu=\frac{xkT}{h}=\frac{2.821439372122\cdot1.38\cdot10^{-23}\cdot1000}{6.62\cdot10^{-34}}\approx58.8\cdot10^{12}Hz=58.8THz"


Infrared radiation (IR).